# Cochain: Balancing Insufficient and Excessive Collaboration in LLM Agent Workflows

### v2

@@ 2.3 Collaboration of LLM Agents
Pangu-1B PMC ${\text{72.28}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.08}}$ ${\text{12.62}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.10}}$ ${\text{22.89}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.14}}$ ${\text{75.67}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.31}}$ ${\text{20.66}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.44}}$ ${\text{29.74}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.26}}$ ${\text{73.96}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.19}}$ ${\text{11.77}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.40}}$ ${\text{19.94}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.43}}$
@@ 4.1 Experimental Setup
We use 17 backbone LLMs, covering both proprietary and open-source models. We compare reasoning-capable models (e.g., DeepSeek-R1-7B DeepSeek-AI (2025)) with standard models (e.g., Pangu-1B Chen et al. (2025), Qwen2-7B 29). We also study using identical backbones across stages versus mixed backbones (Llama2-7B Touvron et al. (2023b), GLM4-9B, Qwen2-7B, DeepSeek-7B) for stage specialization. For open-source backbones, we perform stage-wise fine-tuning on MSCoRe: the agent at stage $k$ is fine-tuned using only the training split of stage-$k$ data, without using any validation/test instances. Proprietary models include GPT-3.5-turbo, GPT-4o Achiam et al. (2023), DeepSeek-V3.1 DeepSeek-AI (2024), and Claude-3.5-haiku.
@@ 4.2 Result
Table 1 summarizes performance across three business domains and four backbones. The results support our central claim that effective workflow collaboration requires balancing Insufficient and Excessive Collaboration. Compared with multi-agent baselines that rely on rich interaction (often leading to excessive collaboration and diluted decisions), Cochain consistently achieves higher quality across domains. Meanwhile, Cochain also improves substantially on a small backbone, indicating that it mitigates insufficient collaboration by injecting stage-aware constraints and dependencies through reusable artifacts rather than relying on stronger backbones alone. For instance, with Pangu-1B on Automotive, Cochain raises BS-F from 72.28 (PMC) to 74.35 and improves ROUGE-L from 22.89 to 26.01, suggesting that the framework helps the agent stay constraint-aware while remaining focused on the core decision.

### v3

@@ 2.3 Collaboration of LLM Agents
openPangu-1B PMC ${\text{72.28}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.08}}$ ${\text{12.62}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.10}}$ ${\text{22.89}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.14}}$ ${\text{75.67}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.31}}$ ${\text{20.66}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.44}}$ ${\text{29.74}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.26}}$ ${\text{73.96}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.19}}$ ${\text{11.77}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.40}}$ ${\text{19.94}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.43}}$
@@
openPangu-7B PMC ${\text{71.41}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.13}}$ ${\text{11.06}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{21.01}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.21}}$ ${\text{70.65}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.41}}$ ${\text{15.77}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.39}}$ ${\text{23.42}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.30}}$ ${\text{71.77}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.23}}$ ${\text{19.55}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.39}}$ ${\text{25.97}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.30}}$
MedAgents ${\text{70.00}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.24}}$ ${\text{10.63}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.28}}$ ${\text{20.26}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.30}}$ ${\text{73.06}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.26}}$ ${\text{18.20}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.21}}$ ${\text{25.35}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.52}}$ ${\text{71.69}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.18}}$ ${\text{19.45}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.36}}$ ${\text{22.89}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.53}}$
Debate(short) ${\text{70.11}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.18}}$ ${\text{10.00}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.21}}$ ${\text{19.78}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.34}}$ ${\text{70.73}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.35}}$ ${\text{12.34}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.26}}$ ${\text{20.77}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.35}}$ ${\text{68.56}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.29}}$ ${\text{18.15}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.23}}$ ${\text{19.62}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.32}}$
Debate(long) ${\text{69.90}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.18}}$ ${\text{9.43}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.21}}$ ${\text{19.56}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.29}}$ ${\text{70.67}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.41}}$ ${\text{13.83}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.22}}$ ${\text{20.73}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.31}}$ ${\text{68.42}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.27}}$ ${\text{17.43}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.36}}$ ${\text{19.44}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.27}}$
CoA ${\text{68.25}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.34}}$ ${\text{7.95}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.35}}$ ${\text{17.32}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.43}}$ ${\text{70.21}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.40}}$ ${\text{14.53}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.26}}$ ${\text{22.98}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.24}}$ ${\text{70.03}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.25}}$ ${\text{19.09}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.29}}$ ${\text{22.80}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.23}}$
Cochain ${\text{74.17}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.08}}$ ${\text{15.49}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.05}}$ ${\text{25.62}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.09}}$ ${\text{70.93}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.09}}$ ${\text{19.40}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.05}}$ ${\text{25.21}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.11}}$ ${\text{73.83}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.07}}$ ${\text{19.60}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.25}}$ ${\text{23.00}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.11}}$
@@ 4.1 Experimental Setup
We use 17 backbone LLMs, covering both proprietary and open-source models. We compare reasoning-capable models (e.g., DeepSeek-R1-7B DeepSeek-AI (2025)) with standard models (e.g., openPangu-7B Chen et al. (2025), Qwen2-7B 29). We also study using identical backbones across stages versus mixed backbones (Llama2-7B Touvron et al. (2023b), GLM4-9B, Qwen2-7B, DeepSeek-7B) for stage specialization. For open-source backbones, we perform stage-wise fine-tuning on MSCoRe: the agent at stage $k$ is fine-tuned using only the training split of stage-$k$ data, without using any validation/test instances. Proprietary models include GPT-3.5-turbo, GPT-4o Achiam et al. (2023), DeepSeek-V3.1 DeepSeek-AI (2024), and Claude-3.5-haiku.
@@ 4.2 Result
Table 1 summarizes performance across three business domains and four backbones. The results support our central claim that effective workflow collaboration requires balancing Insufficient and Excessive Collaboration. Compared with multi-agent baselines that rely on rich interaction (often leading to excessive collaboration and diluted decisions), Cochain consistently achieves higher quality across domains. Meanwhile, Cochain also improves substantially on a small backbone, indicating that it mitigates insufficient collaboration by injecting stage-aware constraints and dependencies through reusable artifacts rather than relying on stronger backbones alone. For instance, with openPangu-1B on Automotive, Cochain raises BS-F from 72.28 (PMC) to 74.35 and improves ROUGE-L from 22.89 to 26.01, suggesting that the framework helps the agent stay constraint-aware while remaining focused on the core decision.